# Can Language Models Reason about

### v1

@@ Abstract
Recent calls for pluralistic alignment emphasize that AI systems should address the diverse needs of all people. Yet, efforts in this space often require sorting people into fixed buckets of pre-specified diversity-defining dimensions (e.g., demographics, personalities, communication styles), risking smoothing out or even stereotyping the rich spectrum of individualistic variations. To achieve an authentic representation of diversity that respects individuality, we propose individualistic alignment. [^1] In this paper, we use the phrase individualistic value to describe “values related to one particular individual,” instead of “values about individualism, such as being independent and self-reliant.” While individualistic alignment can take various forms, in this paper, we introduce [Uncaptioned image] IndieValueCatalog, a dataset transformed from the influential World Values Survey (WVS), to study language models (LMs) on the specific challenge of individualistic value reasoning. Specifically, given a sample of an individual’s value-expressing statements, models are tasked with predicting their value judgments in novel cases. With IndieValueCatalog, we reveal critical limitations in frontier LMs’ abilities to reason about individualistic human values with accuracies only ranging between 55% to 65%. Moreover, our results highlight that a precise description of individualistic values cannot be approximated only via demographic information. We also identify a partiality of LMs in reasoning about global individualistic values, as measured by our proposed Value Inequity Index ($\sigma$Inequity). Finally, we train a series of Individualistic Value Reasoners (IndieValueReasoner) using IndieValueCatalog to enhance models’ individualistic value reasoning capability, revealing new patterns and dynamics into global human values. We outline future research challenges and opportunities for advancing individualistic alignment.
@@ 1 Introduction
Figure 1: Refer to caption IndieValueCatalog, transformed from World Value Survey (WVS), contains statements expressing individualistic human values and preferences from 94K real humans worldwide. With this resource, we study LMs’ ability to reason about individual human values.
@@
Recent advocates for pluralistic alignment (Sorensen et al., 2024; Kirk et al., 2024b) underscore the importance of AI systems being geared towards the diverse perspectives and needs of all people. However, existing methods for achieving this goal (and existing evaluation frameworks for measuring success) face a key limitation—the diversity of people is pre-specified and coarsely categorized. People are often labeled by their cultural, demographic, or community affiliations, papering over the variation of individuals within groups (Feng et al., 2024; Castricato et al., 2024; Sun et al., 2024). Pre-selected diversity-defining dimensions, e.g., demographics (Moon et al., 2024; Kwok et al., 2024), personality (Castricato et al., 2024; Jiang et al., 2023; Serapio-García et al., 2023; Zhu et al., 2024), writing styles (Han et al., 2024; Jang et al., 2023), necessitate sorting individuals into coarse buckets. These choices not only pose the risk of stereotyping (Kirk et al., 2024b), but also inherit potentially negative biases from the specific choice of the diversity dimensions used. While some evaluations exist for assessing value representations among more fine-grained demographic groups (Durmus et al., 2024; Santurkar et al., 2023), these efforts still rely on group-level distributional inferences, and do not directly translate to individual-level variation.
@@
As a bottom-up alternative to addressing these challenges, we propose individualistic value alignment, a maximal version of pluralistic alignment that models diversity at the individual level. This framework focuses on inferring individual preferences from the ground up, bypassing the need for pre-defining categories of people and thereby providing a more authentic representation of diversity by honoring the uniqueness of individuals. As a crucial step towards building individuality-respecting AI, we propose and study individualistic value reasoning—a task for inferring a person’s general value system based on descriptive evidence of their preferences and applying this inference to predict their value preferences in new situations.
@@
One key challenge in studying individual human values lies in the difficulty of acquiring multi-faceted data that is sufficiently representative of an individual’s overall value system. To this end, we present [Uncaptioned image] IndieValueCatalog, a dataset specifically designed to evaluate and advance language models’ ability to reason about an individual’s value preferences in novel situations. IndieValueCatalog transforms unstructured survey questions from the influential social science study of World Value Survey (WVS) into 929 standardized natural language statements describing one’s value preferences (e.g., “I don’t believe in life after death,” “family is not very important in my life”). Our data conversion results in a rich repository of value-expressing statements from 93K unique real humans across the globe. Each person has, on average, 242 and maximally 253 value-expressing statements, along with 31 demographics-declaring statements. In sum, IndieValueCatalog presents the first application of the WVS for studying individualistic human values with LMs in a unified, configurable, and easy-to-measure schema.
@@
With IndieValueCatalog, we first expose the lack of proficiency of frontier LMs in understanding and predicting individualistic human values, as demonstrated by zero-shot accuracies ranging between 55% to 65%. We also introduce Value Inequity Index ($\sigma$Inequity), a unified metric for assessing the degree of equity and impartiality of LMs on this task, which complements metrics measuring LMs’ proficiency in understanding individual values and reveals important shortcomings of LMs. We also discover that adding demographic specifications alongside value-expressing statements has only a marginal impact on improving individualistic value predictions for strong LMs. This highlights the risks of over-relying on demographic factors to define the identities and values of individuals and stresses the importance of addressing values from a granular perspective.
@@
Finally, we train a collection of Individualistic Value Reasoners (IndieValueReasoner) models on IndieValueCatalog, achieving improved proficiency and $\sigma$Inequity on the individualistic value reasoning task, as measured by held-out evaluation data. We conduct extensive experimentation involving different training configurations with IndieValueCatalog, e.g., the number of value-expressing demonstration statements, the granularity of these statements, and the regional origins of the training individuals. Our findings reveal novel dynamics and characteristics of global human values captured in the classical social science resource. We hope our study inspires further research into individualistic value alignment and reasoning, and we outline key challenges and opportunities for future exploration.
@@
Credible, real-world cross-cultural data that captures diverse human values and preferences is difficult to obtain at scale (Castricato et al., 2024). The influential World Value Survey (WVS) addresses this challenge by collecting global responses on social, political, economic, religious, and cultural values (Haerpfer et al., 2020a). With the growing social impact of LMs, WVS data has been used to assess LMs’ biases across demographic groups (Zhao et al., 2024; Durmus et al., 2024). However, for the first time, individual respondent data sequences of WVS are being used to evaluate LMs’ reasoning on individualistic values and preferences.
2.1 Dataset Transformation
Unifying unstructured questions into natural language statements. The original WVS is composed of questions with varying answer formats and fragmented language descriptions. We standardized all multiple-choice and Likert scale questions by converting them into unified natural language statements reflecting value preferences. For instance, we morph questions (e.g., WVS Q131: “Could you tell me how secure you feel these days?”) and answers (e.g., 1. “very secure,” 2. “quite secure” …) into sets of statements like “I feel very secure these days.” Figure 1 and Table 12 show example converted statements in two distinct granularity forms, i.e., polarity-grouped (polar) and refined statements. Demographic questions (31 in total) were similarly converted into identity-declaring statements (e.g., “I’m currently in Andorra”; “I’m an immigrant to this country”). See Table 7-11 for all demographics questions. The full details of data processing are described in Appendix §A.
@@ 2.1 Dataset Transformation
#Questions (Q) #Statements (S-refined) #Statements (S-polar) #Person
@@
Total #Valid Q Avg. #Valid Q per person #Person with full Q set
@@
Table 1: Statistics of IndieValueCatalog data conversion.
Table 1 shows the statistics of IndieValueCatalog. 253 original questions were converted to 929 possible statements for the refined setup and 567 statements for the polar setup, across 93K real humans across the world. For each WVS question, exactly one statement is chosen by each survey respondent (unless a question was omitted by a respondent). The combinatorial answer space for all 253 questions in IndieValueCatalog is extremely large: the refined setup has $1.65\times 10^{139}$ answer combinations and the polar setup has $3.94\times 10^{86}$ combinations, making predicting the exact value choices of a person highly difficult.
2.2 Evaluating LMs on Individualistic Values Reasoning
@@
Table 2: $\sigma$Inequity, i.e., Value Inequity Index, measures the level of partiality or inequity of LMs in reasoning about individualistic human values across diverse population groups averaged by 13 demographic dimensions, e.g., age, income.
@@
Evaluation setups. As illustrated in Figure 1, each individual’s statements are divided into a demonstration (between 50 to 200 statements) and a probing subset (39 statements across 13 WVS question categories; see details in Table 13 of Appendix §B.1 for details of data split). For evaluation, LMs are tasked with selecting the statement most likely to align with the individual’s values from an unseen probing set of value statements based on the demonstration value statements, and optionally, self-declared demographic statements, also from WVS. To facilitate a robust evaluation, we adopt a cross-validation setup with three splits of 200 demonstration questions and 39 probing questions; reporting averaged results to prevent overfitting specific probing set choices. Finally, we sample 800 individuals from IndieValueCatalog as the held-out probing and evaluation set, ensuring a balanced demographic representation.
@@
Formally, $\mathbb{Q}$ is the full set of 253 value-inquiring questions and $\mathbb{I}$ represents all individuals in IndieValueCatalog, which is split into a held-out evaluation subset with 800 individuals ($\mathbb{I}_{\text{eval}}$) and a remaining training subset ($\mathbb{I}_{\text{train}}$). Each question $q\in\mathbb{Q}$ has a set of statements $S_{q}$ expressing varying opinions regarding $q$. For each individual $I_{i}\in\mathbb{I}$, with each question $q\in\mathbb{Q}$, $I_{i}$ chooses one of the statements in $S_{q}$, i.e., $s_{q}^{I_{i}}=S_{q}(I_{i}),s_{q}^{I_{i}}\in S_{q}$, which best represents their opinions regarding $q$. $s_{q}^{I_{i}}$ may be na in cases where the individual does not choose a valid statement option in $S_{q}$.
@@
Each probing setup, $P_{j}\in\{P_{0},P_{1},P_{2}\}$, splits $\mathbb{Q}$ into a probing set of $39$ questions ($\mathbb{Q}_{P_{j}}^{\text{probe}}$) and a remaining demonstration set ($\mathbb{Q}_{P_{j}}^{\text{demo}}$). For each $I_{i}\in\mathbb{I}_{\text{eval}}$ we sample $d$ valid demonstration questions, i.e., $\mathbb{Q}_{P_{j}}^{\text{demo}}(I_{i},d)\subseteq\mathbb{Q}_{P_{j}}^{\text{demo}}$, and gather the chosen statements of $I_{i}$ of these questions, i.e., $\mathbb{S}_{P_{j}}^{\text{demo}}(I_{i},d)=\{s_{q}^{I_{i}}|\forall q\in\mathbb{Q}_{P_{j}}^{\text{demo}}(I_{i},d)\}$. During probing, we present a model, $M$, with $\mathbb{S}_{P_{j}}^{\text{demo}}(I_{i},d)$ along with statement options of all probing questions, $\mathbb{S}_{P_{j}}^{\text{probe}}=\{S_{q}|\forall q\in\mathbb{Q}_{P_{j}}^{\text{probe}}\}$. Finally, we conclude $M$’s choice of value statements for $I_{i}$ of each probing question by sampling from its output, $\{\hat{s}_{M,q}^{I_{i}}\sim M(S_{q}|\mathbb{S}_{P_{j}}^{\text{demo}}(I_{i},d))|\forall q\in\mathbb{Q}_{P_{j}}^{\text{probe}}\}$. We decode with temperature=0 and top_p=1.
@@
Measuring LM’s proficiency in individualistic value reasoning. The average accuracy of $M$ for each individual across all three probing setups and the overall accuracy are calculated as follows.
$Acc_{M}^{I_{i}}=\frac{1}{3\times|\mathbb{Q}_{P_{j}}^{\text{probe}}|}\sum_{P_{j}\in\{P_{0},P_{1},P_{2}\}}\sum_{q\in\mathbb{Q}_{P_{j}}^{\text{probe}}}\mathbbm{1}\big{[}\hat{s}_{M,q}^{I_{i}}=s_{q}^{I_{i}}\big{]}\quad\;$ $\text{and}\quad\;Acc_{M}=\frac{1}{|\mathbb{I}_{\text{eval}}|}\sum_{I_{i}\in\mathbb{I}_{\text{eval}}}Acc_{M}^{I_{i}}$
@@
Measuring LM’s impartiality and equity in individualistic value reasoning. It’s critical to ensure AI development to show an impartially proficient level of understanding of individuals with different demographic characteristics. Here, we introduce Value Inequity Index ($\sigma$Inequity), a metric for measuring the impartiality or equity level of a LM in individualistic value reasoning. In essence, we measure how much performance variance a LM shows in the individualistic value reasoning task across demographic groups—a lower variance means a model shows more impartial understanding across populations. We consider 13 demographic dimensions ($\mathcal{D}^{k}\in\mathbb{D}$; e.g., country of birth, income level, self-assessed social class) from WVS for measuring the cross-group variances (see §B.1 for details). Each demographic dimension is broken into numbers of demographic groups, $g_{k_{t}}\in\mathcal{D}^{k}$; e.g., low/middle/high-income levels for the $\mathcal{D}^{k}$—income level. Every individual belongs to one of the demographic groups for each demographic dimension, i.e., $\mathcal{D}^{k}(I_{i})=g_{k_{t}}^{I_{i}}$. We denote all evaluation individuals who belong to the $g_{k_{t}}$ as $\mathbb{I}_{\text{evel}}^{g_{k_{t}}}=\{I_{i}\;|\;\forall I_{i}\in\mathbb{I}_{\text{evel}},\mathcal{D}^{k}(I_{i})=g_{k_{t}}\}$. We define $\sigma$Inequity of a LM, $M$, as follows.
@@
$\sigma\text{{Inequity}}_{M}=\frac{1}{|\mathbb{D}|}\sum_{\mathcal{D}^{k}\in\mathbb{D}}\sigma(\{Acc_{M}^{\mathbb{I}_{\text{evel}}^{g_{k_{t}}}}\;|\;\forall g_{k_{t}}\in\mathcal{D}^{k}\})$
@@
where $Acc_{M}^{\mathbb{I}_{\text{probe}}^{g_{k_{t}}}}$ is the accuracy among population of the $g_{k_{t}}$ demographic group for model $M$. $\sigma$ denotes standard deviation. Intuitively, $\sigma\text{{Inequity}}_{M}$ represents how much variances the individualistic human value reasoning ability is for $M$ across a range of demographic groups. The lower $\sigma\text{{Inequity}}_{M}$ is, the more impartial $M$ is regarding different demographics groups.
3 Can LMs Reason about Individualistic Human Values?
Figure 3: Evaluation of LMs’ capabilities in reasoning about pluralistic human values and preferences using IndieValueCatalog. Random randomly chooses a statement candidate. GPT-4o (0806) Rand lets GPT-4o randomly guess a statement without demonstration statements.
Model $\sigma$Inequity $\downarrow$
GPT-4o(0806) 3.03
GPT-4o(0513) 2.87
GPT-4o-mini(0718) 2.55
GPT-4-turbo(0409) 2.83
LLama-3.1-8B 2.97
LLama-3.1-70B 1.94
Mixtral-8x7B 3.19
Mixtral-8x22B 3.06
Qwen2-72B 3.24
Claude-3.5(Sonnet) 3.14
Table 3: $\sigma$Inequity, i.e., Value Inequity Index, measures the level of partiality or inequity of LMs in reasoning about individualistic human values across diverse population groups averaged by 13 demographic dimensions, e.g., age, income.
We describe representative probing results below. Please refer to §B.2 for the full experiments.

### v2

@@ Abstract
Recent calls for pluralistic alignment emphasize that AI systems should address the diverse needs of all people. Yet, efforts in this space often require sorting people into fixed buckets of pre-specified diversity-defining dimensions (e.g., demographics), risking smoothing out individualistic variations or even stereotyping. To achieve an authentic representation of diversity that respects individuality, we propose individualistic alignment. [^1] We use the phrase individualistic value to describe “values related to one individual,” instead of “values about individualism, such as being independent and self-reliant.” While individualistic alignment can take various forms, we introduce [Uncaptioned image] IndieValueCatalog, a dataset transformed from the influential World Values Survey (WVS), to study language models (LMs) on the specific challenge of individualistic value reasoning. Given a sample of an individual’s value-expressing statements, models are tasked with predicting this person’s value judgments in novel cases. With IndieValueCatalog, we reveal critical limitations in frontier LMs, which achieve only 55 % to 65% accuracy in predicting individualistic values. Moreover, our results highlight that a precise description of individualistic values cannot be approximated only with demographic information. We also identify a partiality of LMs in reasoning about global individualistic values, as measured by our proposed Value Inequity Index ($\sigma$Inequity). Finally, we train a series of IndieValueReasoners to reveal new patterns and dynamics into global human values.
Can Language Models Reason about
[Uncaptioned image] Individualistic Human Values and Preferences?
Liwei Jiangthanks: Work done while at Allen Institute for Artificial Intelligence.  Taylor Sorensen  Sydney Levine♣∗  Yejin Choi University of Washington  Google DeepMind  Stanford University lwjiang@cs.washington.edu [Uncaptioned image] Code: https://github.com/liweijiang/indievalue
@@ 1 Introduction
Figure 1: Refer to caption IndieValueCatalog contains statements expressing individualistic human values from 94K real humans worldwide. With this resource, we study LMs’ ability to reason about individual human values.
@@
Recent advocates for pluralistic alignment underscore the importance of AI systems being geared towards the diverse perspectives and needs of all people (Sorensen et al., 2024). However, existing methods and evaluation frameworks for achieving this goal face a key limitation—the diversity of people is pre-specified and coarsely categorized, papering over individuality (Castricato et al., 2024; Sun et al., 2024). Pre-selected diversity-defining dimensions, e.g., demographics (Moon et al., 2024; Kwok et al., 2024), personality (Castricato et al., 2024; Jiang et al., 2023; Zhu et al., 2024), cultures (Chiu et al., 2024b), writing styles (Han et al., 2024; Jang et al., 2023), necessitate sorting individuals into coarse buckets. These choices not only pose the risk of stereotyping (Kirk et al., 2024b), but also inherit potentially negative biases from the specific choice of the diversity dimensions. While some evaluations exist for assessing value representations among more fine-grained demographic groups (Durmus et al., 2024; Santurkar et al., 2023), these efforts still rely on group-level distributional inferences, and do not capture individual variations.
@@
As a bottom-up alternative to addressing these challenges, we propose individualistic value alignment, a maximal version of pluralistic alignment that models diversity at the individual level. This framework infers individual preferences from the ground up, avoiding predefined categories and thereby authentically representing diversity by honoring individuality. As a crucial step towards this goal, we propose and study individualistic value reasoning—a task for inferring a person’s general value system based on descriptive evidence of their preferences and applying this inference to predict their value preferences in new situations.
@@
One key challenge in studying individual human values lies in the difficulty of acquiring multi-faceted data that is sufficiently representative of an individual’s overall value preferences. To this end, we present [Uncaptioned image] IndieValueCatalog, a dataset specifically designed to evaluate and advance LMs’ ability to reason about an individual’s value preferences in novel situations. IndieValueCatalog transforms unstructured survey questions from the influential social science study of World Value Survey (WVS) into 929 standardized natural language statements describing one’s value preferences (e.g., “I don’t believe in life after death”). Our data conversion results in a rich repository of value-expressing statements from 93K unique real humans across the globe. Each person has, on average, 242 (max 253) value-expressing statements, along with 31 demographics-declaring statements. In sum, IndieValueCatalog presents the first application of WVS for studying individualistic human values with LMs in a unified, configurable, and easy-to-measure schema.
@@
With IndieValueCatalog, we first expose the lack of proficiency of frontier LMs in interpreting and predicting individualistic human values, as demonstrated by zero-shot accuracies ranging between 55% to 65%. We also introduce Value Inequity Index ($\sigma$Inequity), a unified metric for assessing the degree of equity and impartiality of LMs on this task, revealing critical inequity of LMs in handling individualistic values across population groups. We also discover that adding demographic specifications alongside value-expressing statements has a marginal impact on improving individualistic value predictions for strong LMs. This highlights the risks of over-relying on demographic factors to define the identities and values of individuals and stresses the importance of addressing values from a granular and descriptive perspective.
@@
Finally, we train a collection of Individualistic Value Reasoners (IndieValueReasoner) with IndieValueCatalog, achieving improved proficiency and $\sigma$Inequity on the individualistic value reasoning task, as measured by held-out evaluation data. We conduct extensive experimentation involving different training configurations with IndieValueCatalog, e.g., the number of value-expressing demonstration statements, the granularity of these statements, and the regional origins of the training individuals. Our findings reveal novel dynamics and characteristics of global human values captured in the classical social science resource. We hope our study inspires further research into individualistic value alignment and reasoning, and we outline key challenges and opportunities.
@@
Credible, real-world cross-cultural data that captures diverse human values and preferences is difficult to obtain at scale (Castricato et al., 2024). The influential World Value Survey (WVS) addresses this challenge by collecting global responses on social, political, economic, religious, and cultural values (Haerpfer et al., 2020a), and has been used to assess LMs’ biases across demographic groups (Zhao et al., 2024; Durmus et al., 2024). However, for the first time, individual respondent data sequences of WVS are used to evaluate LMs’ reasoning on individualistic values and preferences.
2.1 Data Transformation
Question Unification.
The original WVS is composed of questions with varying answer formats (e.g., multiple-choice, Likert scale) and fragmented language descriptions. We thus standardized all questions by converting them into unified natural language statements reflecting value preferences. For instance, we morph questions (e.g., WVS Q131: “Could you tell me how secure you feel these days?”) and answers (e.g., 1. “very secure,” 2. “quite secure” …) into sets of statements like “I feel very secure these days.” Figure 1 and Table 11 show example converted statements in two distinct granularity forms, i.e., polarity-grouped (polar) and refined statements. Demographic questions (31 in total) were similarly converted into identity-declaring statements (e.g., “I’m an immigrant to this country”). See Table 6-10 for demographics questions and Appendix §B for full details.
@@ 2.1 Data Transformation
253 original questions are converted to 929 and 567 possible statements for refined and polar setups, respectively, across 93K survey respondents from 70+ countries. For each WVS question, one statement is chosen by each person (unless a question was chosen to be omitted). The combinatorial answer space for all 253 questions is extremely large: the refined setup has $1.65\times 10^{139}$ and the polar setup has $3.94\times 10^{86}$ answer combinations, making predicting the exact value choices of a person highly difficult. Table 1 in Appendix §B shows the dataset statistics.
@@
#Questions (Q) #Refined Stmt #Polar Stmt #Person (P)
@@
Total #Valid Q #Valid Q / P #P w/ Full Q Set
@@
Table 1: Statistics of IndieValueCatalog.
2.2 Evaluation Setups
@@
Table 2: $\sigma$Inequity, i.e., Value Inequity Index, measures the level of partiality or inequity of LMs in reasoning about individualistic human values across diverse population groups averaged by 13 demographic dimensions.
Evaluation setups.
@@
As illustrated in Figure 1, each individual’s statements are divided into a demonstration (50 to 200 statements) and a probing subset (39 statements across 13 WVS question categories; see Table 12 of Appendix §C.1 for data split details). LMs are tasked with selecting a statement most likely to align with an individual’s values among an unseen probing set according to demonstration value statements and, optionally, demographic statements. We adopt a cross-validation setup with three splits of 200 demonstrations and 39 probes; reporting averaged results to prevent overfitting specific probing set choices. Finally, we sample 800 individuals from IndieValueCatalog as the held-out probing and evaluation set, ensuring a balanced demographic representation.
@@
Formally, $\mathbb{Q}$ is the full set of 253 value-inquiring questions and $\mathbb{I}$ represents all individuals in IndieValueCatalog, which is split into a held-out evaluation set with 800 individuals ($\mathbb{I}_{\text{eval}}$) and a training set ($\mathbb{I}_{\text{train}}$). Each question $q\in\mathbb{Q}$ has a set of statements $S_{q}$ expressing varying opinions regarding $q$. For each individual $I_{i}\in\mathbb{I}$, with each question $q\in\mathbb{Q}$, $I_{i}$ chooses one of the statements in $S_{q}$, i.e., $s_{q}^{I_{i}}=S_{q}(I_{i}),s_{q}^{I_{i}}\in S_{q}$, which best represents their opinions regarding $q$.
@@
Each probing setup, $P_{j}\in\{P_{0},P_{1},P_{2}\}$, splits $\mathbb{Q}$ into a probing set of $39$ questions ($\mathbb{Q}_{P_{j}}^{\text{probe}}$) and a remaining demonstration set ($\mathbb{Q}_{P_{j}}^{\text{demo}}$). For each $I_{i}\in\mathbb{I}_{\text{eval}}$ we sample $d$ demonstration questions, i.e., $\mathbb{Q}_{P_{j}}^{\text{demo}}(I_{i},d)\subseteq\mathbb{Q}_{P_{j}}^{\text{demo}}$, and gather the chosen statements of $I_{i}$ of these questions, i.e., $\mathbb{S}_{P_{j}}^{\text{demo}}(I_{i},d)=\{s_{q}^{I_{i}}|\forall q\in\mathbb{Q}_{P_{j}}^{\text{demo}}(I_{i},d)\}$. During probing, we present a model, $M$, with $\mathbb{S}_{P_{j}}^{\text{demo}}(I_{i},d)$ along with statement options of all probing questions, $\mathbb{S}_{P_{j}}^{\text{probe}}=\{S_{q}|\forall q\in\mathbb{Q}_{P_{j}}^{\text{probe}}\}$. Finally, we conclude $M$’s choice of value statements for $I_{i}$ by sampling with temperature=0222We aim to elicit deterministic predictions from the model to avoid randomness in the evaluation resulting from temperature-scaled sampling. and top_p=1, i.e., $\{\hat{s}_{M,q}^{I_{i}}\sim M(S_{q}|\mathbb{S}_{P_{j}}^{\text{demo}}(I_{i},d))|\forall q\in\mathbb{Q}_{P_{j}}^{\text{probe}}\}$.
Measuring proficiency.
@@
The average accuracy of $M$ for each individual across three probing setups and the overall accuracy are calculated as follows.
$Acc_{M}^{I_{i}}=\frac{1}{3\times|\mathbb{Q}_{P_{j}}^{\text{probe}}|}\sum_{P_{j}\in\{P_{0},P_{1},P_{2}\}}\sum_{q\in\mathbb{Q}_{P_{j}}^{\text{probe}}}\mathbbm{1}\big{[}\hat{s}_{M,q}^{I_{i}}=s_{q}^{I_{i}}\big{]}$ (1)
$Acc_{M}=\frac{1}{|\mathbb{I}_{\text{eval}}|}\sum_{I_{i}\in\mathbb{I}_{\text{eval}}}Acc_{M}^{I_{i}}$ (2)
Measuring impartiality and equity.
@@
We introduce Value Inequity Index ($\sigma$Inequity), a metric for measuring the impartiality or equity level of a LM in individualistic value reasoning. In essence, we measure how much performance variance a LM shows in reasoning about individuals across demographic groups—a lower variance means more impartial understanding. We consider 13 demographic dimensions ($\mathcal{D}^{k}\in\mathbb{D}$, e.g., income level; see Appendix §C.1 for details). Each demographic dimension is broken into numbers of groups, $g_{k_{t}}\in\mathcal{D}^{k}$, e.g., low/middle/high for $\mathcal{D}^{\text{income level}}$. Every individual belongs to one of the groups for each demographic dimension, i.e., $\mathcal{D}^{k}(I_{i})=g_{k_{t}}^{I_{i}}$. We denote all individuals belong to $g_{k_{t}}$ as $\mathbb{I}_{\text{evel}}^{g_{k_{t}}}=\{I_{i}\;|\;\forall I_{i}\in\mathbb{I}_{\text{evel}},\mathcal{D}^{k}(I_{i})=g_{k_{t}}\}$. We define $\sigma$Inequity of a LM, $M$, as follows.
@@
$\sigma\text{{Inequity}}_{M}=\frac{1}{|\mathbb{D}|}\sum_{\mathcal{D}^{k}\in\mathbb{D}}\sigma(\{Acc_{M}^{\mathbb{I}_{\text{evel}}^{g_{k_{t}}}}\;|\;\forall g_{k_{t}}\in\mathcal{D}^{k}\})$ (3)
@@
where $Acc_{M}^{\mathbb{I}_{\text{probe}}^{g_{k_{t}}}}$ is the accuracy among population of the $g_{k_{t}}$ demographic group for model $M$. $\sigma$ denotes standard deviation. The lower $\sigma\text{{Inequity}}_{M}$, the more impartial $M$ is.
3 Can LMs Reason about Individualistic Human Values and Preferences?
Proficiency in individualistic value reasoning.